A wheel makes 360 revolutions in one minute. Through how many radians does it turn in one second?

Grade:12th pass

1 Answers

Harshit Singh
3 years ago
Welcome to AskIITians

Given,

Number of revolutions made by the wheel in 1 minute = 360

1 minute = 60 seconds

Number of revolutions in 1 second = 360/60 = 6

Angle made in 1 revolution = 360°

Angles made in 6 revolutions = 6 × 360°

Radian measure of the angle in 6 revolutions = 6 × 360 × π/180

= 6 × 2 × π

= 12π

Hence, the wheel turns 12π radians in one second.

Thanks